5.2.7 A concerned city council member thinks that one of the neighborhoods in the city close to a nuclear power plant has an unusually high rate of cancer. For the whole city, the rate is 1%. Twenty households in that neighborhood are randomly selected, what is the probability that at least 1 household in that neighborhood has an occupant with cancer. Is this event unusual?
Answer #1

P(Atleast 1 occupant with cancer out of 20)

= 1 - P(None occupant with cancer)

= 1 - (1 - 0.01)20

= 1 - (0.99)20

= 0.1821

Since the probability of this event is greater than 0.05, this event is not unusual.